OpenStudy (anonymous):

Is the set {-1,0,1] closed under division?

OpenStudy (unklerhaukus):

can you divide by zero?

OpenStudy (unklerhaukus):

-1 /-1 = 1 -1 / 0 = ????? -1 / 1 = -1 0 /-1 = 0 0 / 0 = ????? 0 / 1 = 0 1 /-1 = -1 1 / 0 = ????? 1 / 1 = 1